OpenStudy (anonymous):

HELP ME UNDERSTAND Solve for x: 3|x − 3| + 2 = 14 No solutions x = −1, x = 8.3 x = 0, x = 7 x = −1, x = 7

OpenStudy (anonymous):

this is english

OpenStudy (anonymous):

wrong subject

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

THE ANSWER IS THE LAST ONE I'LL EXPLAIN FOR YOU

OpenStudy (anonymous):

nevermind if i explain it to you you will be even more confused

OpenStudy (anonymous):

3(|x−3|)+2=14 Step 1: Add -2 to both sides 3(|x−3|)+2+−2=14+−2 3(|x−3|)=12 Step 2: Divide both sides by 3 3(|x−3|)/3 = 12/3 |x−3|=4 Step 3: Solve Absolute Value |x−3|=4 We know either x−3=4 or x−3=−4 x−3=4(Possibility 1) x−3+3=4+3(Add 3 to both sides) x=7 x−3=−4(Possibility 2) x−3+3=−4+3(Add 3 to both sides) x=−1

OpenStudy (anonymous):

Final answer: x=7 or x=−1